The Geometric Mean of 10 items of a series was 16,2
It was later found out that one of the items was wrongly taken as 12.9 instead of 21.9
Calculate correct GM?
kindly solve this riddle
1
Expert's answer
2014-09-24T11:10:03-0400
The formula for geometric mean is

G = (p1*p2*p3...*pn)^(1/n)

where p1, p2,...pn are values of items
If we replace the value p_old (it is among p1, p2,...pn) with another P_new, G will change as

G' =(p_new)^(1/n)* G/((p_old)^(1/n)) =21.9^(1/10)* 16.2/(12.9^(1/10))= 17.08